Log8base4+log2base4
Answer by mathmate(429)


Question:
Evaluate Log8base4+log2base4

Solution:
We will need basic definition of log x to the base b, and some laws of logarithms:
y=logbx means by=x
hence
logbb=1
and
log(a)+log(b)=log(ab)
log(a²)=2log(a)

log4(8)+log4(2)
=log4(8*2)
=log4(16)
=log4(4^2)
=2log4(4)
=2*1
=2